# Thermodynamic & algorithmic efficiency in heterogeneous networks

### v1

@@ Abstract
Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today. At the border between statistical physics, machine learning and network theory, we provide some computational empirical evidence indicating that the emergence of complex structures in a real world environment is the macroscopic thermodynamic result of the optimal trade-off between the ability to encode information about the environment in the very features of the network and the ability to transmit information across it. While the former is quantified by the von Neumann network entropy and the latter by the Helmholtz free energy, the two are gathered into a single quantity tantamount to the thermodynamic efficiency. We take as a case study an algorithm originally proposed for creating and searching trees of patterns under the assumption that, akin to living system, the algorithm seeks the best internal representation of the outer world, namely a real world dataset of patterns. Remarkably, optimal thermodynamic efficiency overlaps with the optimal algorithmic efficiency, defined in terms of accuracy and computational cost, whether the former is measured at the resolution scale in where the network transits from exhibiting a specific heat with a single peak to a scale-invariant region. Moreover the critical hypothesis is investigated and empirical evidence suggests that the system benefits from the critical-state when it confronts with higher-complexity environments, while it does not when the environment is simpler.
@@ 1 Introduction
Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today. We aim at some computational empirical evidence suggesting that the emergence of complex structures is the macroscopic thermodynamic result of optimal ability to store the information about the environment and the ability to transmit information across the structure [1, 2]. The latter is defined in terms of microscopic quantities thanks to the fairly newly developments in statistical physics of fields in network theory [2, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will than encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work a system seeks the best-fitness configuration given some external constraints $r$ and the dataset $\mathcal{D}$.
@@
We take as case study the algorithm introduced by Portegys [11] for nearest-neighbour (1-NN) pattern recognition tasks and subsequently extended to k-NN [12]. The algorithm employs a tree formation protocol where raw data are reorganized into tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing tree structure. At each step the new pattern is linked to either the currently parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the relative distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.  Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star-graph turns out. For some intermediate value of $r$ a ragged tree structure is produced, while for $r\rightarrow 2$ a 1d-lattice is created. Remarkably, the algorithm produces exclusively trees, i.e. number of spanning trees is always 1, which are sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection111 ”Natural selection does not work as an engineer works. It works like a tinkerer - who does not know exactly what he is going to produce but uses whatever he finds around him whether it be pieces of string, fragments of wood, or old cardboards; in short works like a tinkerer who uses everything at his disposal to produce some kind of workable object.”  [13], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a distance based rule.
@@
The Portegys algorithm has the desirable properties of being simple, explainable and subject to a single control parameter $r$. Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning and network science, and we can show that information-theoretical tools as the von Neumann entropy and its derivative reflect the functional capabilities of the network. Most notably thermodynamic efficiency and algorithmic efficiency demonstrate noticeable overlap.
@@ 2.1 Statistical physics of networks
Given an initial state of the network encoding the concentration of quantity, e.g. information, in one or more nodes, $\underline{s}(0)$, the state at time $\tau$ is given by $\underline{s}(\tau)=e^{-\tau\hat{L}}\underline{s}(0)$. Where $\hat{K}=e^{-\tau\hat{L}}$ is the network propagator and is the discrete counter-part of the path integral formulation for general diffusion processes [18, 19].
@@
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [18, 20]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a edge-centric random walker to be back in its initial location at time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
The network entropy thus counts the (logarithmic) number of informationally connected component on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/\log\tau$ it is possible to identify the characteristic resolution scales of the network [20, 3]. For instance, a constant specific heat reflects the scale-invariance of the network where information is lost at a constant rate. Such a situation is typical of d-dimensional lattices and scale invariant networks with power law distributed spectrum. On the contrary, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star-graph, depicted in figure 1, or an Erdös-Rényi network, see purple curve in figure 1. Before the peak the informationally connected components are the single nodes and after the peak the informationally connected components is the full-network reviling a single trivial-scale.
@@
Figure 2: Von Neumann entropy as a function of the diffusion time. Panel (a) depicts diffusion over two exemplary networks with analogous topological features to those depicted in panel (b): in blue a stochastic block model network exhibiting tripartite modular structure, and in purple a Erdös-Rényi random network with the same number of nodes of the blue one. For small $\tau\sim 0$, both networks exhibit the maximum spectral entropy value $\log N$, where $N$ is the number of nodes, reflecting the segregated phase. For $\tau\rightarrow\infty$, the spectral entropy converges to $\log C$, where $C$ is the number of connected components, thus reflecting the integrated phase. For intermediate values of $\tau$, the Erdös-Rényi network shows a single critical scale at which the network transitions from the segregated to the integrated phase. The position on the horizontal axis of the critical scale is related to the edge density of the network as well as the heterogeneity in the degree distribution. On the contrary, the blue network exhibits two critical resolution scales in the entropic transition. At smaller $\tau$, information initially diffuses within the modules ending into a plateau. A later slower diffusion process between the three modules completes the transition to the integrated phase. The insets show the density matrix $\hat{\rho}_{\tau}$ across the entropic transition on the blue curve. The tripartite modular structure of the network is shown by the emergence of the three blocks in the density matrix. Graphs depicted in (b) have only illustration purpose. Illustration from [21].
@@ 2.2 Thermodynamics of networks
The beauty of the statistical physics theory for networks introduced above is that it naturally allows to view network phenomena from a new perspective and to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities as work, $W$, and heat, $Q$, can be defined in (at least mathematical) analogy to their thermodynamic counterparts [2]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network. In turns, lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
@@
On the other hand, it is obvious that adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [2, 4]. Less obvious is how, for fixed number of edges, the degree distribution influences the flow of information. As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [14, 22]. As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter. From $1\leq Z\leq\log N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$. Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
@@
which is tantamount to the work in the canonical ensemble. Moreover, $U=W+Q\geq 0$ which implies $W\geq|Q|$. A formation process can be depicted as the trade-off between this two necessities of a networked complex systems: the maximization of information flow, while at the same time maintening an amount of entropy such that the network can encode and express sufficient amount of diversity trough its traits.
@@ 3 Results
Our results are evaluated over subsets of three different datasets: the NIST [23], the MNIST [24] and the FashionMNIST [25] datasets. These data sets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure. The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep- learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. In all our experiments a subset of 2000 samples is extracted from the train-set and the test-set independently. In the Suppplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashioMNIST. The 10 classes are equally represented both in the test and in training set. We limit to $r\leq 0.85$ as the accuracy, and hence the algoritmic efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
3.1 Algorithmic efficiency
After the tree formation described in section 1, the Portegys algorithm enters the search phase, where the best match to a test pattern is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the performance would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure. The search process is an iterative process which begins with the computation of the distance between the test pattern and the root-node pattern scaled by $r$. The distance is then compared to the distance between the test pattern and all of the root node’s children. The first node with a pattern that has a smaller distance is than selected as the new root node, and the process starts over with the newly selected node as root. The search phase terminates when no node whose pattern has smaller distance to the test pattern than the current one can be found.
It is natural to define an algorithmic efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
$\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
the logarithm is taken as the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes. For the case of a star-graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of figure 3.
Figure 3: Algorithmic efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best-match, (b). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency. The shaded grey area designates the standard deviation. The purple shaded area in panel (c) indicates the intersection between the blue curve and the standard deviation at the optimal point whose y-axis projection is indicated by yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST dataset. The 10 digit-classes are equally represented both in the train- and test-set.
3.2 Thermodynamic and algorithmic efficiency at the emergence of new C-peaks
In figures 4, 6 and 5 methodology and results showing the coherence between algorithmic and thermodynamic efficiency are replicated for the three investigated datasets. For each dataset, panel (a) depicts the $\tau$-resolution scales identified by the first peak (blue curve), $\tau^{(1)}_{peak}$, and last peak (green curve), $\tau^{(2)}_{peak}$, of the specific heat, $C$, as a function over the control parameter $r$. See figure S5 in the Supplementary for more von Neumann entropy and specific heat curves as a function of $r$.

### v2

# Thermodynamic & pattern-matching efficiency in heterogeneous networks

@@ Abstract
Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today. At the border between statistical physics, machine learning and network theory, we provide some computational empirical evidence indicating that the emergence of complex structures in the outer world is the macroscopic thermodynamic result of the optimal trade-off between the ability to encode information about the environment in the very features of the network and the ability to transmit information across it. While the former is quantified by the von Neumann network entropy and the latter by the Helmholtz free energy, the two are combined into a single quantity tantamount to the thermodynamic efficiency. We take as a case study an algorithm originally proposed for creating and searching trees of patterns under the assumption that, akin to living system, the algorithm seeks the best internal representation of the outer world, namely a real world dataset of patterns. Remarkably, the optimal thermodynamic efficiency overlaps with the optimal pattern-matching efficiency, defined in terms of accuracy and computational cost, whether the former is measured at the resolution scale where the network transits from exhibiting a specific heat with a single peak to a scale-invariant region. Moreover the critical hypothesis is investigated and empirical evidence suggests that the system benefits from the vicinity to the critical point whether it is faced with higher-complexity environments, while it does not when the environment is simpler.
@@ 1 Introduction
Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today. We aim at some computational empirical evidence suggesting that the emergence of complex structures is the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [1, 2]. The latter is defined in terms of microscopic quantities thanks to the fairly newly developments in statistical physics of fields in network theory [2, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will than encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
@@
We take as case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbour (1-NN) pattern recognition tasks and subsequently extended to k-NN [13]. The algorithm employs a tree formation protocol where raw data are reorganized into tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing tree structure. At each step the new pattern is linked to either the currently parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the relative distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.  Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out. For some intermediate value of $r$ a ragged tree structure is produced, while for $r\rightarrow 2$ a 1d-lattice is created. Remarkably, the algorithm produces exclusively trees, i.e. number of spanning trees is always 1, which are sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection111 ”Natural selection does not work as an engineer works. It works like a tinkerer - who does not know exactly what he is going to produce but uses whatever he finds around him whether it be pieces of string, fragments of wood, or old cardboards; in short works like a tinkerer who uses everything at his disposal to produce some kind of workable object.”  [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a distance based rule.
@@
The Portegys algorithm has the desirable properties of being simple, explainable and subject to a single control parameter $r$. Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network. Most notably thermodynamic efficiency and pattern-matching efficiency demonstrate noticeable overlap.
@@ 2.1 Statistical physics of networks
Given an initial state of the network encoding the concentration of quantity, e.g. information, in one or more nodes, $\underline{s}(0)$, the state at time $\tau$ is given by $\underline{s}(\tau)=e^{-\tau\hat{L}}\underline{s}(0)$. Where $\hat{K}=e^{-\tau\hat{L}}$ is the network propagator and is the discrete counterpart of the path integral formulation for general diffusion processes [19, 20].
@@
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [19, 21]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location at time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/\log\tau$ it is possible to identify the characteristic resolution scales of the network [21, 3]. For instance, a constant specific heat reflects the scale-invariance of the network where information is lost at a constant rate. Such a situation is typical of d-dimensional lattices and scale invariant networks with power law distributed spectrum. On the contrary, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network, see purple curve in Figure 1. Before the peak the informationally connected components are the single nodes and after the peak the informationally connected components components are the full-network revealing a single trivial-scale.
@@
Figure 2: Von Neumann entropy as a function of the diffusion time. Panel (a) depicts diffusion over two exemplary networks with analogous topological features to those depicted in panel (b): in blue a stochastic block model network exhibiting tripartite modular structure, and in purple a Erdös-Rényi random network with the same number of nodes of the blue one. For small $\tau\sim 0$, both networks exhibit the maximum spectral entropy value $\log N$, where $N$ is the number of nodes, reflecting the segregated phase. For $\tau\rightarrow\infty$, the spectral entropy converges to $\log C$, where $C$ is the number of connected components, thus reflecting the integrated phase. For intermediate values of $\tau$, the Erdös-Rényi network shows a single critical scale at which the network transitions from the segregated to the integrated phase. The position on the horizontal axis of the critical scale is related to the edge density of the network as well as the heterogeneity in the degree distribution. On the contrary, the blue network exhibits two critical resolution scales in the entropic transition. At smaller $\tau$, information initially diffuses within the modules ending into a plateau. A later slower diffusion process between the three modules completes the transition to the integrated phase. The insets show the density matrix $\hat{\rho}_{\tau}$ across the entropic transition on the blue curve. The tripartite modular structure of the network is shown by the emergence of the three blocks in the density matrix. Graphs depicted in (b) have only illustration purpose. Illustration from [22, 21, 23].
@@ 2.2 Thermodynamics of networks
The statistical physics theory for networks introduced above naturally allows to view network phenomena from a new perspective and to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities as work, $W$, and heat, $Q$, can be defined in (at least mathematical) analogy to their thermodynamic counterparts [2]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network. In turns, lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
@@
On the other hand, it is obvious that adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [2, 4]. Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information. As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [15, 24]. As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter. From $1\leq Z\leq N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$. Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
@@
which is tantamount to the work in the canonical ensemble. Moreover, $U=W+Q\geq 0$ which implies $W\geq|Q|$. A formation process can be depicted as the trade-off between these two necessities of the networked complex system: the need to maximize the information flow, while at the same time retaining an amount of structural entropy such that the network can encode and express sufficient diversity through its traits.
@@
3 Methods
Our results are evaluated over subsets of three different datasets: the NIST [25], the MNIST [26] and the FashionMNIST [27] datasets. These datasets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure. The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep- learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. In all our experiments a subset of 2000 samples is extracted from the train-set and the test-set independently. In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST. The 10 classes are equally represented both in the test and in training set. We limit to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$. The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] and where distance between patterns is the Euclidean distance.
3.1 Pattern-matching efficiency
After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to a test pattern is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure. The search process is an iterative process which begins with the computation of the distance between the test pattern and the root-node pattern scaled by $r$. The distance is then compared to the distance between the test pattern and all of the root node’s children. The first node with a pattern that has a smaller distance is than selected as the new root node, and the process starts over with the newly selected node as root. The search phase terminates when no node whose pattern has smaller distance to the test pattern than the current one can be found.
It is natural to define an pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
$\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes. For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of Figure 3.
Figure 3: Pattern-matching efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best match, (b). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency. The shaded grey area designates the standard deviation. The purple shaded area in panel (c) indicates a tolerance $r$-domain of optimal efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST dataset. The 10 digit-classes are equally represented both in the train- and test-set.
@@ 4 Results
4.1 Thermodynamic and pattern-matching efficiency at the emergence of new C-peaks
Figure 4 shows the coherence between algorithmic defined in Section 3 and thermodynamic efficiency in Equation 5 for the three investigated datasets. In panels (a)-(c) the pattern-matching efficiency, $\theta$, is depicted. The vertical red-dashed line indicates the $r$ value for the optimal algorithmic trade-off, while the shaded purple region indicates a $r$-domain tolerance of the optimal pattern-matching efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off. In panels (d)-(f) the thermodynamic efficiency, $\eta$, shows remarkable overlap with the pattern-matching efficiency $\theta$ in terms of optimal $r$. See the $r$-range spanned by the purple-shaded areas in panels (a)-(c) and the vertical dashed lines in panels (d)-(f). The overlap between the two efficiencies suggests that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles. See the next Section for the discussion on the double peak of the $\eta$ curves for the MNIST.